OpenStudy (anonymous):

Write the equation of the line parallel to the line 10x - 5y = 8 and passing through the point (2,4).

OpenStudy (gorv):

for parrallel slop will be same so m=2 euation of parallel line will be y-4=2(x-2)

OpenStudy (gorv):

10x-5y=8 10x-8=5y y=2x-8/5 on comparing with y=mx+c slope=2

OpenStudy (gorv):

got it